OpenStudy (anonymous):

Solve 3x^2 + 5x - 2 = 0 by factoring. a) 1/3, -2 b) 1, -3 c) 2, -12 d) -1/3, -2

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x^2 + 6 x - x - 2 = 0 }\) helps?

OpenStudy (anonymous):

Um not really lol. I'm not smart when it comes to math

OpenStudy (anonymous):

these is another way to write 5x=6x-x \[3x(x+2)-1(x+2)=(3x-1)(x+2)\] grouping

OpenStudy (anonymous):

So it would be c?

OpenStudy (anonymous):

a. (x + 2) (3 x - 1) = 0 (x + 2) = 0, x = -2 (3 x - 1) = 0, x = 1/3

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Thank you for answering

OpenStudy (anonymous):

You are welcome.